# Non-toric brane webs, Calabi–Yau 3-folds, and 5d SCFTs

### v1

@@ Abstract.
We study webs of 5-branes with 7-branes in Type IIB string theory from a geometric perspective. To any such a web $W$, we attach a log Calabi–Yau surface $(Y,D)$ with a line bundle $L$. We then describe supersymmetric webs, which are webs defining 5d superconformal field theories (SCFTs), in terms of the geometry of $(Y,D,L)$. We also introduce particular supersymmetric webs called “consistent webs”, and show that any 5d SCFT defined by a supersymmetric web can be obtained from a consistent web by adding free hypermultiplets. Using birational geometry of degenerations of log Calabi–Yau surfaces, we provide an algorithm to test the consistency of a web in terms of its dual polygon. Moreover, for a consistent web $W$, we provide an algebro-geometric construction of the mirror $\mathcal{X}^{\mathrm{can}}$ to $(Y,D,L)$, as a non-toric canonical 3-fold singularity, and show that M-theory on $\mathcal{X}^{\mathrm{can}}$ engineers the same 5d SCFT as $W$. We also explain how to derive explicit equations for $\mathcal{X}^{\mathrm{can}}$ using scattering diagrams, encoding disk worldsheet instantons in the A-model, or equivalently the BPS states of an auxiliary rank one 4d $\mathcal{N}=2$ theory.
@@ 1.1. Background and context
One of the most remarkable predictions of string/M-theory is the existence of 5-dimensional superconformal field theories (5d SCFTs) [90]. Constructing such SCFTs has been of significant interest, and two powerful approaches for doing so are provided by M-theory on canonical $3$-fold singularities on one hand [35, 61, 83, 96], and intersecting branes in Type IIB string theory on the other [2]. A natural question is to compare these two approaches. The aim of this paper is to address this question: given a 5d SCFT engineered by a configuration of branes in Type IIB string theory, can we obtain the same 5d SCFT from M-theory on a dual canonical 3-fold singularity? Conversely, can we give an algebro-geometric description of the class of canonical 3-fold singularities that admit a dual Type IIB string description in terms of intersecting branes?
The answer to these questions is well-known when the configuration of branes in Type IIB string theory is just a web of 5-branes [2, 75]. In this situation, the dual canonical 3-fold singularity is a toric affine Calabi–Yau 3-fold, which has an explicit combinatorial description, obtained from the web of 5-branes. Moreover, every toric affine Calabi–Yau 3-fold arises this way. However, answering these questions is significantly more challenging, when considering $5$-branes together with $7$-branes [18, 32]. In this situation, while the existence of non-toric dual canonical 3-fold singularities is still predicted via string dualities [18], giving a precise algebro-geometric construction of such singularities has been a longstanding open question – see [13, 22] for recent progress in this direction. The present paper provides a complete answer to this question, building on recent mathematical advances on mirror symmetry for log Calabi–Yau surfaces.
@@ 1.2.1. Webs of 5-branes and toric mirror symmetry
In §2, we review the known construction of the M-theory dual toric canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ to an asymptotic web of 5-branes $\overline{W}^{\mathrm{asym}}$ [2, 75]. In particular, we explain how to view $\overline{\mathcal{X}}^{\mathrm{can}}$ as the mirror to the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ with momentum polytope $\overline{P}$ dual to $\overline{W}^{\mathrm{asym}}$, where $\overline{Y}$ is a projective toric surface, $\overline{D}$ is the toric boundary divisor, and $\overline{L}$ is an ample line bundle on $\overline{Y}$. Mirrors to crepant resolutions of $\overline{\mathcal{X}}^{\mathrm{can}}$ are then given by maximal polarized toric degenerations of $(\overline{Y},\overline{D},\overline{L})$ associated to webs $\overline{W}$ with asymptotic web $\overline{W}^{\mathrm{asym}}$ – see Figure 2.1 . In this paper, we provide a natural generalization of this mirror construction to the case of webs of 5-branes with 7-branes.
@@ 2. Webs of 5-branes and toric mirror symmetry
2.1. Webs of $5$-branes, $5$d SCFTs, and dual polygons
2.1.1. Webs of $5$-branes and $5$d SCFTs
@@
[…]
Figure 2.1. A web of $5$-branes $\overline{W}$ on the left, and the associated asymptotic web of $5$-branes $\overline{W}^{\mathrm{asym}}$ on the right.
Starting with any web of 5-branes $\overline{W}$, we obtain an associated asymptotic web $\overline{W}^{\mathrm{asym}}$, by first scaling $\overline{W}$ by $t>0$ and then taking the limit as $t\rightarrow 0$ – this amounts to contracting all edges on $\overline{W}$, as illustrated in Figure 2.1. The webs $\overline{W}$ with associated asymptotic web $\overline{W}^{\mathrm{asym}}$ can be viewed as deformations of $\overline{W}^{\mathrm{asym}}$. From the viewpoint of the 5d SCFT, deforming the web of branes from $\overline{W}^{\mathrm{asym}}$ to $\overline{W}$ amounts to deforming the theory either by mass parameters or by moving on the Coulomb branch of the theory – see [28] for an exposition. The union of the Coulomb branches of all the massive deformations of the 5d SCFT can be described as the moduli space of all webs $\overline{W}$ with asymptotic web $\overline{W}^{\mathrm{asym}}$ and considered up to translations of $\mathbb{R}^{2}$.
@@ 2.1.2. The dual polygon to an asymptotic web of $5$-branes
Refer to caption$(\overline{P},\overline{\mathcal{P}})$$\overline{P}$
Figure 2.2. The polygon $\overline{P}$ dual to the web $\overline{W}^{\mathrm{asym}}$, with polyhedral decomposition $\overline{\mathcal{P}}$ dual to the web $\overline{W}$ in Figure 2.1
Throughout this paper, we will assume that we are not in the degenerate situation where $\overline{W}^{\mathrm{asym}}$ is contained in a line and $\overline{P}$ is a line segment – elsewise there won’t be any intersecting branes, thus no non-trivial $5$d SCFT.
Webs of 5-branes $\overline{W}$ with associated asymptotic web $\overline{W}^{\mathrm{asym}}$ are dual to integral polyhedral decompositions $\mathcal{\overline{P}}$ of the polygon $\overline{P}$ dual to $\overline{W}^{\mathrm{asym}}$, which are regular in the sense that their faces are the domains of linearity of a continuous piecewise linear function on $\overline{P}$. We will often denote by $(\overline{P},\overline{\mathcal{P}})$ a lattice polygon $\overline{P}$ endowed with a regular polyhedral decomposition $\overline{\mathcal{P}}$. For example, webs of 5-branes that are obtained as generic deformations of $\overline{W}$ are 3-valent and are dual to regular integral triangulations of $\overline{P}$ into lattice triangles of size one, that is, with three sides of integral length one. Note that a lattice triangle of size one is integral-affinely equivalent to the standard simplex in $\mathbb{Z}^{2}$ with vertices $(0,0)$, $(1,0)$, and $(0,1)$.
@@ 2.2.1. Asymptotic webs of $5$-branes and Calabi–Yau $3$-folds
The 3-fold $\overline{Z}$ is an affine conic bundle over $(\mathbb{C}^{\star})^{2}$, which is degenerate over the algebraic curve $\overline{C}^{\circ}$, as illustrated in Figure 2.3.
[…]
Figure 2.3. The Calabi–Yau 3-fold $\overline{Z}$ on the left, and the compactification $\overline{C}$ of $\overline{C}^{\circ}$ in $Y$ on the right.
@@ 2.2.2. Asymptotic webs of $5$-branes and polarized toric surfaces
The monomials $x^{a}y^{b}$ with $(a,b)\in\overline{P}_{\mathbb{Z}}$ naturally define a basis of sections of $\overline{L}$, and so the curve $\overline{C}^{\circ}\subset(\mathbb{C}^{\star})^{2}$ admits a natural compactification $\overline{C}\subset\overline{Y}$ in the linear system of $\overline{L}$. If the coefficients $c_{a,b}$ of $f$ as in Equation (2.1) are all non-zero, then one can show that the curve $\overline{C}$ does not pass through the 00-dimensional strata of $\overline{D}$ – see Figure 2.3.
@@ 2.2.3. Webs of 5-branes and degenerations of polarized toric surfaces
Let $\overline{\mathcal{P}}$ be a regular polyhedral decomposition of $\overline{P}$, $\overline{\varphi}$ an integral piecewise-linear function on $\overline{P}$ defining $\overline{\mathcal{P}}$, and $\overline{W}$ the corresponding integral dual web of 5-branes. Following [85], we associate to this data a toric degeneration of the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ into unions of polarized toric varieties glued along toric strata. We sketch the construction below.
Consider the polyhedral decomposition $\overline{\mathcal{P}}^{\vee}$ of $\mathbb{R}^{2}$ defined by $\overline{W}$, as in [85, §3]. Then, the union of the convex polyhedral cones, indexed by the cells $\Xi$ of $\overline{\mathcal{P}}^{\vee}$ and given by the closures in $\mathbb{R}^{2}\times\mathbb{R}$ of the cones $\mathbb{R}_{\geq 0}\cdot(\Xi\times\{1\})$, form a $3$-dimensional toric fan $\Sigma$. In addition, the projection map to the last coordinate defines a map of fans $\Sigma\to\Sigma_{\mathbb{C}}=\mathbb{R}_{\geq 0}$ to the toric fan of $\mathbb{C}$, and so a toric morphism
@@
The general fiber $\overline{\mathcal{Y}}_{t}$ and the central fiber $\overline{\mathcal{Y}}_{0}$ can be described by examining the preimages of the cones $\{0\}$ and $\mathbb{R}_{\geq 0}$ respectively, under the map of fans $\Sigma\to\Sigma_{\mathbb{C}}$ – see [85, Lemma $3.4$] and [85, Proposition $3.5$ ] for details. It follows by construction that the preimage of the cone $\{0\}$ of $\Sigma_{\mathbb{C}}$ is the asymptotic web $\overline{W}^{\mathrm{asym}}$ associated to $\overline{W}$ (see [85, pg 9]), and hence the general fiber is the corresponding toric surface $\overline{Y}$. Moreover, the preimage of any point in the interior of $\mathbb{R}_{\geq 0}=\Sigma_{\mathbb{C}}$ is $\mathbb{R}^{2}$ with the polyhedral decomposition $\overline{\mathcal{P}}^{\vee}$, which therefore can be viewed as the dual intersection complex of the central fiber $\overline{\mathcal{Y}}_{0}$. In particular, corresponding to each vertex $v$ of $\overline{\mathcal{P}}^{\vee}$, there is a toric irreducible component $(\overline{Y}_{0}^{v},\partial\overline{Y}_{0}^{v})$ of the central fiber $\overline{\mathcal{Y}}_{0}$, where we denote by $\partial\overline{Y}_{0}^{v}$ the toric boundary divisor.
By standard toric geometry, there exists a unique line bundle $\overline{\mathcal{L}}$ on $\mathcal{Y}$ such that, for every toric curve $C_{e}\simeq\mathbb{P}^{1}$ in $\mathcal{Y}$ corresponding to an edge or leg $e$ of $\overline{W}$, we have $\mathrm{deg}\,\overline{\mathcal{L}}|_{C_{e}}=w_{e}$, where $w_{e}$ is the weight of $e$ in $\overline{W}$. The restriction of $\overline{\mathcal{L}}$ to a general fiber $(\overline{Y},\overline{D})$ is isomorphic to the ample line bundle $\overline{L}$. Moreover, $(\overline{P},\overline{\mathcal{P}})$ is the intersection complex of the polarized central fiber, as in [52]. In particular, for every vertex $v$ of $\overline{\mathcal{P}}^{\vee}$, the momentum polygon of the polarized toric surface $(\overline{Y}_{0}^{v},\partial{\overline{Y}}_{0}^{v},\overline{\mathcal{L}}|_{\overline{Y}_{0}^{v}})$ is the face of $\overline{\mathcal{P}}$ corresponding to $v$.
@@ 3.1.1. Asymptotic configurations of $7$-branes
The point $x_{i}$ in Definition 3.1 is viewed as the position in $\mathbb{R}^{2}$ of a $(p_{i},q_{i})$ 7-brane, see Figure 3.1. An asymptotic configuration of 7-branes defines an integral affine manifold with singularities $B$ obtained from $\mathbb{R}^{2}$ by cutting along the half-lines $x_{i}+\mathbb{R}_{\geq 0}(p_{i},q_{i})$ and imposing that the clockwise monodromy around $x_{i}$ through the cut is given by the matrix $M_{(p_{i},q_{i})}$. The integral affine manifold with singularities $B$ is well-defined even if $(p_{i},q_{i})=(p_{j},q_{j})$ for $i\neq j$, since the matrices $M_{(p_{i},q_{i})}$ and $M_{(p_{j},q_{j})}$ are equal, and so in particular commute.
[…]
Figure 3.1. An asymptotic configuration of 7-branes.
@@
A Hanany–Witten move along a $(p_{i},q_{i})$ $7$-brane $x_{i}$ is the operation of bringing the position of $x_{i}$ across the origin along the monodromy-invariant direction $\mathbb{R}(p_{i},q_{i})$, so that $x_{i}$ is now contained in the ray $\mathbb{R}_{<0}(p_{i},q_{i})$. In particular, this new configuration of 7-branes is no longer an asymptotic configuration of 7-branes as in Definition 3.1. However, it can be transformed into one by viewing $x_{i}$ as a $(-p_{i},-q_{i})$ 7-brane, and rotating the cut in the clockwise direction from $x_{i}+\mathbb{R}_{\geq 0}(p_{i},q_{i})$ to $x_{i}+\mathbb{R}_{<0}(p_{i},q_{i})$. During such a rotation, the cut of the $(p_{i},q_{i})$ 7-brane intersects the $(p_{j},q_{j})$ 7-branes with $\det((p_{i},q_{i}),(p_{j},q_{j}))<0$. Such a $(p_{j},q_{j})$ 7-brane crosses the cut of the $(p_{i},q_{i})$ 7-brane in the anti-clockwise direction around $x_{i}$, and so becomes a $(p_{j}^{\prime},q_{j}^{\prime})$ 7-brane, where

### v2

@@ Abstract.
We study webs of 5-branes with 7-branes in Type IIB string theory from a geometric perspective. Mathematically, a web of 5-branes with 7-branes is a tropical curve in $\mathbb{R}^{2}$ with focus-focus singularities introduced. To any such a web $W$, we attach a log Calabi–Yau surface $(Y,D)$ with a line bundle $L$. We then describe supersymmetric webs, which are webs defining 5d superconformal field theories (SCFTs), in terms of the geometry of $(Y,D,L)$. We also introduce particular supersymmetric webs called “consistent webs”, and show that any 5d SCFT defined by a supersymmetric web can be obtained from a consistent web by adding free hypermultiplets. Using birational geometry of degenerations of log Calabi–Yau surfaces, we provide an algorithm to test the consistency of a web in terms of its dual polygon. Moreover, for a consistent web $W$, we provide an algebro-geometric construction of the mirror $\mathcal{X}^{\mathrm{can}}$ to $(Y,D,L)$, as a non-toric canonical 3-fold singularity, and show that M-theory on $\mathcal{X}^{\mathrm{can}}$ engineers the same 5d SCFT as $W$. We also explain how to derive explicit equations for $\mathcal{X}^{\mathrm{can}}$ using scattering diagrams, encoding disk worldsheet instantons in the A-model, or equivalently the BPS states of an auxiliary rank one 4d $\mathcal{N}=2$ theory.
@@ 1.1. Background and context
One of the most remarkable predictions of string/M-theory is the existence of 5-dimensional superconformal field theories (5d SCFTs) [93]. Constructing such SCFTs has been of significant interest, and two powerful approaches for doing so are provided by M-theory on canonical $3$-fold singularities on one hand [9, 10, 11, 38, 64, 86, 99], and intersecting branes in Type IIB string theory on the other [2]. A natural question is to compare these two approaches. The aim of this paper is to address this question: given a 5d SCFT engineered by a configuration of branes in Type IIB string theory, can we obtain the same 5d SCFT from M-theory on a dual canonical 3-fold singularity? Conversely, can we give an algebro-geometric description of the class of canonical 3-fold singularities that admit a dual Type IIB string description in terms of intersecting branes?
The answer to these questions is well-known when the configuration of branes in Type IIB string theory is just a web of 5-branes, mathematically described by a tropical curve in $\mathbb{R}^{2}$ [2, 78]. In that setting, the dual canonical 3-fold singularity is a Gorenstein toric affine Calabi–Yau 3-fold, which has an explicit combinatorial description, obtained from the web of 5-branes. Moreover, every Gorenstein toric affine canonical 3-fold singularity arises this way. However, answering these questions is significantly more challenging, when considering $5$-branes together with $7$-branes, mathematically described by a tropical curve in $\mathbb{R}^{2}$ with focus-focus singularities introduced [21, 35]. In this situation, while the existence of non-toric dual canonical 3-fold singularities is still predicted via string dualities [21], giving a precise algebro-geometric construction of such singularities has been a longstanding open question – see [16, 25] for recent progress in this direction. The present paper provides a complete answer to this question, building on recent mathematical advances on mirror symmetry for log Calabi–Yau surfaces.
@@ 1.2.1. Webs of 5-branes and toric mirror symmetry
In §2, we review the known construction of the M-theory dual toric canonical 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ to an asymptotic web of 5-branes $\overline{W}^{\mathrm{asym}}$ [2, 78]. In particular, we explain how to view $\overline{\mathcal{X}}^{\mathrm{can}}$ as the mirror to the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ with momentum polytope $\overline{P}$ dual to $\overline{W}^{\mathrm{asym}}$, where $\overline{Y}$ is a projective toric surface, $\overline{D}$ is the toric boundary divisor, and $\overline{L}$ is an ample line bundle on $\overline{Y}$. Mirrors to crepant resolutions of $\overline{\mathcal{X}}^{\mathrm{can}}$ are then given by maximal polarized toric degenerations of $(\overline{Y},\overline{D},\overline{L})$ associated to webs $\overline{W}$ with asymptotic web $\overline{W}^{\mathrm{asym}}$ – see Figure 1. In this paper, we provide a natural generalization of this mirror construction to the case of webs of 5-branes with 7-branes.
@@ 2. Webs of 5-branes and toric mirror symmetry
2.1. Webs of $5$-branes, 55d SCFTs, and dual polygons
2.1.1. Webs of $5$-branes and 55d SCFTs
@@
Figure 1. A web of $5$-branes $\overline{W}$ on the left, and the associated asymptotic web of $5$-branes $\overline{W}^{\mathrm{asym}}$ on the right.
Starting with any web of 5-branes $\overline{W}$, we obtain an associated asymptotic web $\overline{W}^{\mathrm{asym}}$, by first scaling $\overline{W}$ by $t>0$ and then taking the limit as $t\rightarrow 0$ – this amounts to contracting all edges on $\overline{W}$, as illustrated in Figure 1. The webs $\overline{W}$ with associated asymptotic web $\overline{W}^{\mathrm{asym}}$ can be viewed as deformations of $\overline{W}^{\mathrm{asym}}$. From the viewpoint of the 5d SCFT, deforming the web of branes from $\overline{W}^{\mathrm{asym}}$ to $\overline{W}$ amounts to deforming the theory either by mass parameters or by moving on the Coulomb branch of the theory – see [31] for an exposition. The union of the Coulomb branches of all the massive deformations of the 5d SCFT can be described as the moduli space of all webs $\overline{W}$ with asymptotic web $\overline{W}^{\mathrm{asym}}$ and considered up to translations of $\mathbb{R}^{2}$.
@@ 2.1.2. The dual polygon to an asymptotic web of $5$-branes
Figure 2. The polygon $\overline{P}$ dual to the web $\overline{W}^{\mathrm{asym}}$, with polyhedral decomposition $\overline{\mathcal{P}}$ dual to the web $\overline{W}$ in Figure 1
Throughout this paper, we will assume that we are not in the degenerate situation where $\overline{W}^{\mathrm{asym}}$ is contained in a line and $\overline{P}$ is a line segment – elsewise there won’t be any intersecting branes, thus no non-trivial 55d SCFT.
Webs of 5-branes $\overline{W}$ with associated asymptotic web $\overline{W}^{\mathrm{asym}}$ are dual to integral polyhedral decompositions $\mathcal{\overline{P}}$ of the polygon $\overline{P}$ dual to $\overline{W}^{\mathrm{asym}}$, which are regular in the sense that their faces are the domains of linearity of a convex continuous piecewise linear function on $\overline{P}$. We will often denote by $(\overline{P},\overline{\mathcal{P}})$ a lattice polygon $\overline{P}$ endowed with a regular polyhedral decomposition $\overline{\mathcal{P}}$. For example, webs of 5-branes that are obtained as generic deformations of $\overline{W}$ are 3-valent and are dual to regular integral triangulations of $\overline{P}$ into lattice triangles of size one, that is, with three sides of integral length one. Note that a lattice triangle of size one is integral-affinely equivalent to the standard simplex in $\mathbb{Z}^{2}$ with vertices $(0,0)$, $(1,0)$, and $(0,1)$.
@@ 2.2.1. Asymptotic webs of $5$-branes and Calabi–Yau $3$-folds
The 3-fold $\overline{Z}$ is an affine conic bundle over $(\mathbb{C}^{\star})^{2}$, which is degenerate over the algebraic curve $\overline{C}^{\circ}$, as illustrated in Figure 3.
Figure 3. The Calabi–Yau 3-fold $\overline{Z}$ on the left, and the compactification $\overline{C}$ of $\overline{C}^{\circ}$ in $Y$ on the right.
@@ 2.2.2. Asymptotic webs of $5$-branes and polarized toric surfaces
The monomials $x^{a}y^{b}$ with $(a,b)\in\overline{P}_{\mathbb{Z}}$ naturally define a basis of sections of $\overline{L}$, and so the curve $\overline{C}^{\circ}\subset(\mathbb{C}^{\star})^{2}$ admits a natural compactification $\overline{C}\subset\overline{Y}$ in the linear system of $\overline{L}$. If the coefficients $c_{a,b}$ of $f$ as in Equation (2.1) are all non-zero, then the curve $\overline{C}$ does not pass through the 0-dimensional strata of $\overline{D}$ – see Figure 3.
@@ 2.2.3. Webs of 5-branes and degenerations of polarized toric surfaces
Let $\overline{\mathcal{P}}$ be a regular polyhedral decomposition of $\overline{P}$ and $\overline{\varphi}$ a convex continuous piecewise-linear function on $\overline{P}$ defining $\overline{\mathcal{P}}$ and having integral values on the integral points of $\overline{P}$. Let $\overline{W}$ be the corresponding dual web of 5-branes. Following [88], we associate to this data a toric degeneration of the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ into unions of polarized toric varieties glued along toric strata. We sketch the construction below.
Consider the polyhedral decomposition $\overline{\mathcal{P}}^{\vee}$ of $\mathbb{R}^{2}$ defined by $\overline{W}$, as in [88, §3]. Then, the union of the convex polyhedral cones, indexed by the cells $\Xi$ of $\overline{\mathcal{P}}^{\vee}$ and given by the closures in $\mathbb{R}^{2}\times\mathbb{R}$ of the cones $\mathbb{R}_{\geq 0}\cdot(\Xi\times\{1\})$, form a $3$-dimensional toric fan $\Sigma$. In addition, the projection map to the last coordinate defines a map of fans $\Sigma\to\Sigma_{\mathbb{C}}=\mathbb{R}_{\geq 0}$ to the toric fan of $\mathbb{C}$, and so a toric morphism
@@
The general fiber $\overline{\mathcal{Y}}_{t}$ and the central fiber $\overline{\mathcal{Y}}_{0}$ can be described by examining the preimages of the cones $\{0\}$ and $\mathbb{R}_{\geq 0}$ respectively, under the map of fans $\Sigma\to\Sigma_{\mathbb{C}}$ – see [88, Lemma $3.4$] and [88, Proposition $3.5$ ] for details. It follows by construction that the preimage of the cone $\{0\}$ of $\Sigma_{\mathbb{C}}$ is the asymptotic web $\overline{W}^{\mathrm{asym}}$ associated to $\overline{W}$ (see [88, pg 9]), and hence the general fiber is the corresponding toric surface $\overline{Y}$. Moreover, the preimage of any point in the interior of $\mathbb{R}_{\geq 0}=\Sigma_{\mathbb{C}}$ is $\mathbb{R}^{2}$ with the polyhedral decomposition $\overline{\mathcal{P}}^{\vee}$, which therefore can be viewed as the dual intersection complex of the central fiber $\overline{\mathcal{Y}}_{0}$. In particular, corresponding to each vertex $v$ of $\overline{\mathcal{P}}^{\vee}$, there is a toric irreducible component $(\overline{Y}_{0}^{v},\partial\overline{Y}_{0}^{v})$ of the central fiber $\overline{\mathcal{Y}}_{0}$, where we denote by $\partial\overline{Y}_{0}^{v}$ the toric boundary divisor.
By standard toric geometry, there exists a unique line bundle $\overline{\mathcal{L}}$ on $\overline{\mathcal{Y}}$ such that, for every toric curve $C_{e}\simeq\mathbb{P}^{1}$ in $\overline{\mathcal{Y}}$ corresponding to an edge or leg $e$ of $\overline{W}$, we have $\mathrm{deg}\,\overline{\mathcal{L}}|_{C_{e}}=w_{e}$, where $w_{e}$ is the weight of $e$ in $\overline{W}$. The restriction of $\overline{\mathcal{L}}$ to a general fiber $(\overline{Y},\overline{D})$ is isomorphic to the ample line bundle $\overline{L}$. Moreover, $(\overline{P},\overline{\mathcal{P}})$ is the intersection complex of the polarized central fiber, as in [55]. In particular, for every vertex $v$ of $\overline{\mathcal{P}}^{\vee}$, the momentum polygon of the polarized toric surface $(\overline{Y}_{0}^{v},\partial{\overline{Y}}_{0}^{v},\overline{\mathcal{L}}|_{\overline{Y}_{0}^{v}})$ is the face of $\overline{\mathcal{P}}$ corresponding to $v$.
@@ 3.1.1. Asymptotic configurations of $7$-branes
The point $x_{i}$ in Definition 3.1 is viewed as the position in $\mathbb{R}^{2}$ of a $(p_{i},q_{i})$ 7-brane, see Figure 4. An asymptotic configuration of 7-branes defines an integral affine manifold with singularities $B$ obtained from $\mathbb{R}^{2}$ by cutting along the half-lines $x_{i}+\mathbb{R}_{\geq 0}(p_{i},q_{i})$ and imposing that the clockwise monodromy around $x_{i}$ through the cut is given by the matrix $M_{(p_{i},q_{i})}$. The integral affine manifold with singularities $B$ is well-defined even if $(p_{i},q_{i})=(p_{j},q_{j})$ for $i\neq j$, since the matrices $M_{(p_{i},q_{i})}$ and $M_{(p_{j},q_{j})}$ are equal, and so in particular commute.
Figure 4. An asymptotic configuration of 7-branes.
@@
A Hanany–Witten move along a $(p_{i},q_{i})$ 77-brane $x_{i}$ is the operation of bringing the position of $x_{i}$ across the origin along the monodromy-invariant direction $\mathbb{R}(p_{i},q_{i})$, so that $x_{i}$ is now contained in the ray $\mathbb{R}_{<0}(p_{i},q_{i})$. In particular, this new configuration of 7-branes is no longer an asymptotic configuration of 7-branes as in Definition 3.1. However, it can be transformed into one by viewing $x_{i}$ as a $(-p_{i},-q_{i})$ 7-brane, and rotating the cut in the clockwise direction from $x_{i}+\mathbb{R}_{\geq 0}(p_{i},q_{i})$ to $x_{i}+\mathbb{R}_{<0}(p_{i},q_{i})$. During such a rotation, the cut of the $(p_{i},q_{i})$ 7-brane intersects the $(p_{j},q_{j})$ 7-branes with $\det((p_{i},q_{i}),(p_{j},q_{j}))<0$. Such a $(p_{j},q_{j})$ 7-brane crosses the cut of the $(p_{i},q_{i})$ 7-brane in the anti-clockwise direction around $x_{i}$, and so becomes a $(p_{j}^{\prime},q_{j}^{\prime})$ 7-brane, where